OpenStudy (anonymous):

log base 5 (x-4)=0 how do i solve this?

OpenStudy (amistre64):

log(1) = 0

OpenStudy (anonymous):

can you explain how you got that?

OpenStudy (amistre64):

if need be, base 5 each side x-4 = 1

OpenStudy (anonymous):

Convert it into exponential form

OpenStudy (amistre64):

its a property of logs ....

OpenStudy (anonymous):

X=5, by converting 5^0=x-4 as 5^0is 1

OpenStudy (amistre64):

let there be a function such that: L(ab) = L(a) + L(b) define a*1 L(a*1) = L(a) + L(1) L(a) = L(a) + L(1) L(a) - L(a) = L(1) 0 = L(1)

OpenStudy (anonymous):

Got it! thank you so much!